OpenStudy (anonymous):

how do I write 0.0462-0.369i in polar form using the exponential form

OpenStudy (anonymous):

mostly a calculator exercise

OpenStudy (anonymous):

\[ 0.0462-0.369i =r\left(\cos(\theta)+i\sin(\theta)\right)\] \[r=\sqrt{a^2+b^2}=\sqrt{(0.0462)^2+(0.369)^2}\]

OpenStudy (anonymous):

and \[\tan(\theta)=\frac{b}{a}=-\frac{0.369}{.0462}\] making \[\theta=\tan^{-1}(-\frac{0.369}{.0462})\]